Give the equation of the horizontal asymptote of the rational function.
g(x) = (5)/(x+6)
these are my options. i think it should be x=0 but it dose not give me that, so y=0??
A) y = 5
B) y = 0
C) y = 5/6
D) x = -6

Answer by RAY100(1637)

Horizontal asymptotes,,, general form,,,y= (a(n)x^n +...)/(b(m)x^m +.....)
.
if n less than m,,,,y=0,,,,as is this case
.
if n=m,,,,,y= a(n) / b(m)
.
if n>m,,,,,no horizontal asymptote
.
in this case y=5x^0/ (x+6),,,n=o,,,,m=1,,n less than m,,,,y=0 is horiz asymptote